# Refer to caption Refer to caption

### v1

@@ Abstract.
scene generation, video diffusion, neural rendering
*Equal contribution. Project Page: https://the-world-explorer.github.io
ccs: Computing methodologies Image-based renderingccs: Computing methodologies Volumetric models
@@ 1. Introduction
One of the core challenges in generating 3D environments is enabling users to freely explore them in an interactive fashion. That is, to create immersive surroundings, users must be able to navigate the scene and look around objects, all while maintaining realistic view synthesis. One line of existing methods leverages $\text{360}^{\circ}$ panoramas and converts them into 3D scenes (Yang et al., 2024b; Zhou et al., 2024). Despite high-quality view synthesis near the scene’s center, their ability to support broader exploration is very limited; occlusions and missing multi-view information result in distorted geometry and inconsistent appearance under larger viewpoint changes. Alternative approaches generate scenes through iterative render-refine-repeat pipelines (Shriram et al., 2024; Höllein et al., 2023; Yu et al., 2024a), often relying on T2I models and monocular depth estimators (Yang et al., 2024a). While this enables them to produce entire scenes beyond panoramas, their iterative nature often leads to stretching or distortion artifacts, since the object geometry is expanded frame-by-frame. More recently, methods using video diffusion models have demonstrated locally plausible object geometry (Sun et al., 2024; Chen et al., 2025b; Ren et al., 2025), i.e., their world prior over continuous motions helps to mitigate such artifacts. However, their generated environments cannot yet be interactively explored, since they either do not generate entire 360 degree scenes or still suffer from distortion artifacts under novel view far beyond centered perspectives.
@@
To address these challenges, we propose a novel scene generation method leveraging camera-guided video generation, that enables free exploration of text-generated 3D environments (Figure 1). Our approach combines the strengths of panorama-based initialization, video diffusion models, and iterative scene generation. First, we construct a $360^{\circ}$ panorama from multiple images, describing separate areas of a diverse environment. By utilizing depth estimation (Yang et al., 2024a) and T2I inpainting (Labs, 2023), we create an initial scene scaffold. Next, we leverage video diffusion models in an iterative generation scheme to explore and expand the areas described by these initial observations. Concretely, we generate multiple small videos, that move into different scene parts and thus provide more multi-view information about their contained objects. We define the video trajectories a-priori, i.e., a diverse set of directions and perspectives iteratively explores the entire $360^{\circ}$ scene. To ensure view-consistency between the separate videos, we propose a scene memory mechanism. It builds a conditional input for the video model from our scene scaffold and the most suitable views among all previously generated images. We further develop a collision detection mechanism, that dynamically adjusts the trajectory length based on the actual generated content, i.e., we avoid degenerate outputs like moving the camera into walls or objects. Finally, all generated views are fused into a single 3D scene representation via 3D Gaussian Splatting optimization (Kerbl et al., 2023). We initialize this reconstruction by estimating a point cloud from all generated images using  (Wang et al., 2025a) and rigidly align it with our generated cameras. This enables interactive exploration of generated 3D worlds from arbitrary viewpoints in real-time based on rasterization.
@@
  • We propose an iterative scene expansion strategy using video diffusion models, driven by trajectory sampling and adaptive collision detection.
  • We design a scene memory mechanism that conditions each video generation step on relevant past frames, improving view consistency and overall scene coherence.
@@ 2. Related Work
2.1. Image- and Video-Generation Models
@@
Text-to-image (T2I) models can create high-quality, diverse images from text (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022). Their visual quality continuously improved over recent years (Podell et al., 2023; Labs, 2023; Xie et al., 2024; Zhang et al., 2023c). As they are trained on billions of images (Schuhmann et al., 2022), they represent strong 2D world priors, that can be adopted for many downstream tasks like inpainting/editing (Brooks et al., 2023; Lugmayr et al., 2022), semantic controls (Zhang et al., 2023a; Ye et al., 2023; Mou et al., 2024), or personalization (Ruiz et al., 2023; Hu et al., 2022). We leverage T2I models in the panorama initialization stage of our method to create large and diverse environments.
Recently, text-to-video (T2V) models can generate worlds under continuously changing perspectives (Blattmann et al., 2023; Yang et al., 2024c; Seawead et al., 2025). A popular line of T2V generators models the diffusion process (Ho et al., 2020) using flow-matching (Lipman et al., 2022) with a U-Net (Ronneberger et al., 2015) or diffusion transformer (DiT) (Peebles and Xie, 2023; Vaswani et al., 2017) architecture. Inspired by the impressive quality and diversity, we leverage these models to build entire 3D scenes autoregressively.
2.2. 3D Scene Generation
@@
Several different approaches exist, that leverage T2I/T2V models as priors for 3D generation. Score-distillation-sampling (Poole et al., 2023) optimizes for diverse objects (Siddiqui et al., 2024; Chen et al., 2023; Wang et al., 2023) or scenes (Cohen-Bar et al., 2023; Zhang et al., 2024; Li et al., 2024b). Others create multi-view consistent images and turn those into 3D scenes (Szymanowicz et al., 2025; Höllein et al., 2024; Gao et al., 2024). In order to create entire indoor/outdoor scenes, (Höllein et al., 2023; Schult et al., 2024; Yu et al., 2024b; Shriram et al., 2024) utilize monocular depth-estimators (Yang et al., 2024a; Ke et al., 2024) in an iterative generation pipeline (Liu et al., 2021). Alternatively, (Wang et al., 2024; Schwarz et al., 2025; Li et al., 2024a) turn $\text{360}^{\circ}$ panoramas into 3D Gaussian Splatting (Kerbl et al., 2023) scenes. Recently, camera-guided video diffusion models are utilized in similar iterative pipelines (Sun et al., 2024; Yu et al., 2024c; Ren et al., 2025). We leverage video models in a novel iterative pipeline to generate large-scale 3D scenes that allow interactive exploration.
2.3. Camera-Controlled Video Generation
@@
Having control over the camera trajectory of generated videos is a desirable property for world generation. Recent approaches finetune T2V models with additional pose input (Bahmani et al., 2024b, a; Liang et al., 2024; He et al., 2025; Bai et al., 2025). Others render point clouds into novel views to condition the next-frame generation (Liu et al., 2024; Wang et al., 2025b; Zhang et al., 2025). While the motion of the videos can be precisely controlled, their length is often limited (e.g., small rotations around single objects). By using T2V models in a sliding-window fashion, (Song et al., 2025; Zhou et al., 2025) generate longer and consistent videos. However, to create entire 3D scenes, it is crucial to generate consistent outputs even when revisiting previously generated areas or when viewing them from different perspectives. To this end, we exploit camera-controlled T2V models in a novel iterative scheme, i.e., backed by a scene memory mechanism, we generate multiple small videos in a 3D-consistent fashion.
@@
Figure 2. Method overview. We generate a fully navigable 3D scene from text input in three stages. Left: we create four initial images with a T2I model (Labs, 2023), and we define that they look outwards, distributed uniformly around the scene center (without overlap). We leverage monocular depth estimators (Yang et al., 2024a) and inpainting to create the remaining four images of our scene scaffold. Mid: we propose a novel iterative video trajectory generation pipeline. It exploits camera-guided video diffusion models (Zhou et al., 2025) to generate additional images of the scene along pre-defined trajectories. For each video generation, we select optimal conditioning views from all previously generated frames, which encourages 3D-consistent content generation. We further discard frames (marked as grey), that are too close to (running into) objects to avoid degenerate outputs. We repeat this trajectory generation multiple times to obtain a final set of posed images. Right: we optimize a 3D Gaussian-Splatting scene (Kerbl et al., 2023) from our generated images. To this end, we first obtain a point-cloud initialization using (Wang et al., 2025a). Then we globally align the predicted points with our camera poses using rigid transformations. Finally, we can explore the scene, i.e., synthesize novel views from perspectives beyond the scene center, in real-time using standard computer graphics rasterization.
@@ 3.1. Review of Video Diffusion Models
[…]
@@ 3.3. Iterative Video Trajectory Generation
Figure 3. Conditional video generation via scene memory. We generate videos in multiple batches of 21 frames along 4 trajectories (blue, purple, red, orange). These trajectories start from each of our initial panoramic images (bottom left). To ensure 3D-consistent outputs, we prefill the first 8 frames with our initial scene scaffold (marked in green). The remaining 5 are sampled based on rotational similarity from our scene memory, which stores the images and poses of all previously generated frames (top). After each trajectory is generated, we append all frames starting from the third image to our scene memory for conditioning of later frames.
Given the initial scene scaffold $\{\mathbf{I}_{i}\}_{i=1}^{8}$, we create additional observations with the video diffusion model. Concretely, we propose an iterative generation scheme: starting from an initial image, we utilize four pre-defined trajectories $\{\mathbf{\pi}^{\text{tgt}}_{k}\}_{k=1}^{4}$ that explore the scene by moving “into and around” that area. In other words, we generate the videos $\mathbf{\hat{V}}_{4i+k}{\sim}p_{\theta}(\mathbf{I}^{\text{tgt}}|\mathbf{I}^{\text{src}},\mathbf{\pi}^{\text{src}},\mathbf{\pi}^{\text{tgt}}_{k})$ for all $i,k$. We iterate this principle multiple times to create in total 32 videos (4 per initial image). Next, we explain the generation in more detail.
@@
Camera-guided video diffusion models such as  (Zhou et al., 2025) are powerful priors for generating locally plausible multi-view observations (e.g., rotating around objects). In contrast, generating entire 3D scenes along a single, large camera path remains challenging due to several reasons. First, defining a continuous camera path a-priori, that covers an entire 3D scene is ambiguous. In other words, we have to adapt the trajectory on-the-fly based on the generated content (e.g., prevent running into walls, resolve occlusions, discover novel generated areas). Second, to achieve 3D-consistency in large 3D worlds, the model must memorize previously generated objects/structures across a large history of previous frames. While approaches like  (Zhou et al., 2025; Song et al., 2025) show promising results, their models still suffer from inconsistencies, collisions, and catastrophic forgetting, when re-visiting previously observed areas (also see Figure 5 and the supplementary material).
@@
Crucially, resolving collisions becomes a much easier task this way: since we already split scene generation into discontinuous and small camera paths, we only need to prevent such degenerate cases for each trajectory separately. In other words, the next trajectory is not impacted by a collision of previous trajectories, which avoids dynamically adjusting our camera paths. Concretely, we first generate the entire video and then utilize a video depth estimator (Chen et al., 2025a) to predict normalized depth maps $\mathbf{\hat{D}}\in[0,1]^{N\times H\times W}$ for all generated frames. We define a fixed area of interest as a binary mask $\mathbf{M}\in\{0,1\}^{H\times W}$, where pixels $(u,v)$ within a center crop of extent $(w,h)$ are set to 1:
@@
Our method leverages the two-stage video generation approach of (Zhou et al., 2025), where anchor frames are first generated uniformly along the trajectory, followed by interpolation between these anchors. Our scene memory mechanism focuses on ensuring consistency of the anchor frames across trajectories. The subsequent interpolation then naturally produces scene-consistent intermediate frames.

### v2

@@ Abstract.
scene generation, video diffusion, 3D Gaussian Splatting
journalyear: 2025copyright: ccconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763946isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Renderingccs: Computing methodologies Volumetric modelsccs: Computing methodologies Reconstruction
@@ 1. Introduction
One of the core challenges in generating 3D environments is enabling users to freely explore them in an interactive fashion. That is, to create immersive surroundings, users must be able to navigate the scene and look around objects, all while maintaining realistic view synthesis. One line of existing methods leverages $\text{360}^{\circ}$ panoramas and converts them into 3D scenes (Yang et al., 2024b; Zhou et al., 2024). Despite high-quality view synthesis near the scene’s center, their ability to support broader exploration is very limited; occlusions and missing multi-view information result in distorted geometry and inconsistent appearance under larger viewpoint changes. Alternative approaches generate scenes through iterative render-refine-repeat pipelines (Shriram et al., 2024; Höllein et al., 2023; Yu et al., 2024a), often relying on T2I models and monocular depth estimators (Yang et al., 2024a). While this enables them to produce entire scenes beyond panoramas, their iterative nature often leads to stretching or distortion artifacts, since the object geometry is expanded frame-by-frame. More recently, methods using video diffusion models have demonstrated locally plausible object geometry (Sun et al., 2024; Chen et al., 2025b; Ren et al., 2025), i.e., their world prior over continuous motions helps to mitigate such artifacts. However, the generated worlds cannot yet be interactively explored, as they either lack full $360^{\circ}$ scene coverage or exhibit distortion artifacts under novel views beyond centered perspectives.
@@
To address these challenges, we propose a novel scene generation method leveraging camera-guided video generation, that enables free exploration of text-generated 3D environments (Figure 1). Our approach combines the strengths of panorama-based initialization, video diffusion models, and iterative scene generation. First, we construct a $360^{\circ}$ panorama from multiple images, describing separate areas of a diverse environment. By utilizing depth estimation (Yang et al., 2024a) and T2I inpainting (Labs, 2023), we create this initial scene scaffold. Next, we leverage video diffusion models in an iterative generation scheme to explore and expand the areas described by these initial observations. Concretely, we generate multiple small videos, that move into different scene parts and thus provide more multi-view information about their contained objects. We define the video trajectories a-priori, i.e., a diverse set of directions and perspectives iteratively explores the entire $360^{\circ}$ scene. To ensure view-consistency between the separate videos, we propose a scene memory mechanism. It builds a conditional input for the video model from our scene scaffold and the most suitable views among all previously generated images. We further develop a collision detection mechanism, that dynamically adjusts the trajectory length based on the actual generated content, i.e., we avoid degenerate outputs like moving the camera into walls or objects. Finally, all generated views are fused into a single 3D scene representation via 3D Gaussian Splatting optimization (Kerbl et al., 2023). We initialize this reconstruction by estimating a point cloud from all generated images using  (Wang et al., 2025a) and rigidly align it with our generated cameras. This enables interactive exploration of generated 3D worlds from arbitrary viewpoints in real-time based on rasterization.
@@
  • We propose an autoregressive scene expansion via video diffusion, driven by trajectory sampling and adaptive collision detection.
  • We design a scene memory to condition each generation step on key prior frames, ensuring view consistency and scene coherence.
@@ 2. Related Work
Image- and Video-Generation Models
@@
Text-to-image (T2I) models can create high-quality, diverse images from text (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022). Their visual quality continuously improved (Podell et al., 2023; Labs, 2023; Xie et al., 2024; Zhang et al., 2023b). As they are trained on billions of images (Schuhmann et al., 2022), they represent strong 2D world priors, that can be adopted for many downstream tasks like inpainting/editing (Brooks et al., 2023; Lugmayr et al., 2022), semantic controls (Zhang et al., 2023a; Ye et al., 2023; Mou et al., 2024), or personalization (Ruiz et al., 2023; Hu et al., 2022). We leverage T2I models in our panorama initialization stage to create large and diverse environments. Video models can generate worlds under continuously changing perspectives (Blattmann et al., 2023; Yang et al., 2024c; Seawead et al., 2025). A popular approach trains diffusion models (Ho et al., 2020) via flow-matching (Lipman et al., 2022) with U-Net (Ronneberger et al., 2015) or DiT (Peebles and Xie, 2023; Vaswani et al., 2017) backbones. We leverage these models to build entire 3D scenes autoregressively.
3D Scene Generation
@@
Several approaches use 2D models as priors for 3D generation. Score distillation sampling (Poole et al., 2023) optimizes for diverse objects (Siddiqui et al., 2024; Chen et al., 2023; Wang et al., 2023) or scenes (Cohen-Bar et al., 2023; Zhang et al., 2024; Li et al., 2024b). Others create multi-view consistent images for 3D reconstruction (Szymanowicz et al., 2025; Höllein et al., 2024; Gao et al., 2024). In order to create entire scenes, (Höllein et al., 2023; Schult et al., 2024; Yu et al., 2024b; Shriram et al., 2024) utilize depth estimation (Yang et al., 2024a; Ke et al., 2024) and inpainting in an iterative generation pipeline (Liu et al., 2021). Alternatively, (Wang et al., 2024; Schwarz et al., 2025; Li et al., 2024a) turn $\text{360}^{\circ}$ panoramas into 3D Gaussian Splatting (Kerbl et al., 2023) scenes. Recently, camera-guided video diffusion models are utilized in similar pipelines (Sun et al., 2024; Yu et al., 2024c; Ren et al., 2025; Liu et al., 2024b). We leverage such video models in a novel iterative pipeline to generate large-scale 3D scenes that allow interactive exploration.
Camera-Controlled Video Generation
@@
Camera control in video generation is beneficial for world generation. Recent approaches finetune models with additional pose input (Bahmani et al., 2024, 2025; Liang et al., 2024; He et al., 2025; Bai et al., 2025). Others render point clouds into novel views to condition the next frame generation (Liu et al., 2024a; Wang et al., 2025b; Zhang et al., 2025). While the motion of the videos can be precisely controlled, their length is often limited (e.g., small rotations around objects). By using T2V models in a sliding-window fashion, (Song et al., 2025; Zhou et al., 2025) generate longer and consistent videos. However, to create entire 3D scenes, it is crucial to generate consistent outputs even when revisiting previously generated areas or when viewing them from different perspectives. To this end, we exploit these models in a novel iterative scheme. Backed by a scene memory mechanism, we generate multiple small videos in a 3D-consistent fashion.
@@
Figure 2. Method overview. We generate a fully navigable 3D scene from text input in three stages. Left: we create four initial images with a T2I model (Labs, 2023), and we define that they look outwards, distributed uniformly around the scene center (without overlap). We leverage monocular depth estimators (Yang et al., 2024a) and inpainting to create the remaining four images of our scene scaffold. Mid: we propose a novel iterative video trajectory generation pipeline. It exploits camera-guided video diffusion models (Zhou et al., 2025) to generate additional images of the scene along pre-defined trajectories. For each video generation, we select optimal conditioning views from all previously generated frames, which encourages 3D-consistent content generation. We further discard frames (marked as grey), that are too close to (running into) objects to avoid degenerate outputs. We repeat this trajectory generation multiple times to obtain a final set of posed images. Right: we optimize a 3D Gaussian-Splatting scene (Kerbl et al., 2023) from our generated images. To this end, we first obtain a point-cloud initialization using (Wang et al., 2025a). Then we align the predicted points to our camera poses via rigid transformation and scaling. Finally, we can explore the scene, i.e., synthesize novel views from perspectives beyond the scene center, in real-time using standard computer graphics rasterization.
[…]
Method overview. We generate a fully navigable 3D scene from text input in three stages. Left: we create four initial images with a T2I model (Labs, 2023), and we define that they look outwards, distributed uniformly around the scene center (without overlap). We leverage monocular depth estimators (Yang et al., 2024a) and inpainting to create the remaining four images of our scene scaffold. Mid: we propose a novel iterative video trajectory generation pipeline. It exploits camera-guided video diffusion models (Zhou et al., 2025) to generate additional images of the scene along pre-defined trajectories. For each video generation, we select optimal conditioning views from all previously generated frames, which encourages 3D-consistent content generation. We further discard frames (marked as grey), that are too close to (running into) objects to avoid degenerate outputs. We repeat this trajectory generation multiple times to obtain a final set of posed images. Right: we optimize a 3D Gaussian-Splatting scene (Kerbl et al., 2023) from our generated images. To this end, we first obtain a point-cloud initialization using (Wang et al., 2025a). Then we align the predicted points to our camera poses via rigid transformation and scaling. Finally, we can explore the scene, i.e., synthesize novel views from perspectives beyond the scene center, in real-time using standard computer graphics rasterization.
@@ 3.1. Review of Video Diffusion Models
Video diffusion (Blattmann et al., 2023; Yang et al., 2024c; Seawead et al., 2025) models the conditional probability distribution $p_{\theta}(\mathbf{I}^{\text{tgt}}|\mathbf{I}^{\text{src}})$ over data $(\mathbf{I}^{\text{tgt}},\mathbf{I}^{\text{src}}){\sim}q(\mathbf{V})$, where $\mathbf{V}{=}\{\mathbf{I}_{1},\mathbf{I}_{2},...,\mathbf{I}_{N}\}$ is a dataset of continuous videos with $\mathbf{I}{\in}\mathbb{R}^{3\times H\times W}$ being the individual images. To achieve more control over the generated video samples, a camera condition $(\mathbf{\pi}^{\text{src}},\mathbf{\pi}^{\text{tgt}})$ is added to each image, which typically is encoded into dense Plücker embeddings (Zhou et al., 2025; Liang et al., 2024). The network $\theta$ then parameterizes the conditional diffusion denoising probability $p_{\theta}(\mathbf{I}^{\text{tgt}}|\mathbf{c}){=}\int p_{\phi}(\mathbf{I}^{\text{tgt}}_{0{:}T}|\mathbf{c})d\mathbf{I}^{\text{tgt}}_{1{:}T}$ with $\mathbf{c}{=}(\mathbf{I}^{\text{src}},\mathbf{\pi}^{\text{src}},\mathbf{\pi}^{\text{tgt}})$ and we progressively add more Gaussian noise to the latent variables $\mathbf{I}^{\text{tgt}}_{1{:}T}$ following (Ho et al., 2020; Lipman et al., 2022). We leverage the recent camera-guided video diffusion model from (Zhou et al., 2025) as a strong prior over the real world. In other words, we exploit the model to generate entire 3D scenes autoregressively, i.e., we generate a large set of images expressed as multiple shorter videos $\{\mathbf{\hat{V}_{1}},...,\mathbf{\hat{V}_{M}}\}$. To ensure high-quality and 3D-consistent video generation, we develop a novel conditioning mechanism that selects the optimal $\mathbf{c}_{k}$ from all previously generated data $\{\mathbf{\hat{V}_{1}},...,\mathbf{\hat{V}_{k-1}}\}$ (see Section 3.3).
@@ 3.3. Iterative Video Trajectory Generation
Figure 3. Conditional video generation via scene memory. We sequentially generate videos in multiple batches of 21 frames along 4 trajectories (blue, purple, red, orange). These trajectories start from each of our initial panoramic images (bottom left). To ensure 3D-consistent outputs, we prefill the first 8 frames with our initial scene scaffold (marked in green). The remaining 5 are sampled based on rotational similarity from our scene memory, which stores the images and poses of all previously generated frames (top). After each trajectory is generated, we append all frames starting from the third image to our scene memory for conditioning of later frames.
[…]
Conditional video generation via scene memory. We sequentially generate videos in multiple batches of 21 frames along 4 trajectories (blue, purple, red, orange). These trajectories start from each of our initial panoramic images (bottom left). To ensure 3D-consistent outputs, we prefill the first 8 frames with our initial scene scaffold (marked in green). The remaining 5 are sampled based on rotational similarity from our scene memory, which stores the images and poses of all previously generated frames (top). After each trajectory is generated, we append all frames starting from the third image to our scene memory for conditioning of later frames.
Given the initial scene scaffold $\{\mathbf{I}_{i}\}_{i=1}^{8}$, we create additional observations with the video diffusion model. Concretely, we propose an iterative generation scheme: starting from an initial image, we utilize four pre-defined trajectories $\{\mathbf{\pi}^{\text{tgt}}_{k}\}_{k=1}^{4}$ that explore the scene by moving “into and around” that area. In other words, we sequentially generate the videos $\mathbf{\hat{V}}_{4i+k}{\sim}p_{\theta}(\mathbf{I}^{\text{tgt}}|\mathbf{I}^{\text{src}},\mathbf{\pi}^{\text{src}},\mathbf{\pi}^{\text{tgt}}_{k})$ for all $i,k$, where suitable condition frames $\mathbf{I}^{\text{src}}$ are selected per video from all previous frames. We iterate this principle multiple times to create in total 32 videos. Next, we explain the generation in more detail.
@@
Camera-guided video diffusion models such as  (Zhou et al., 2025) are powerful priors for generating locally plausible multi-view observations (e.g., rotating around objects). In contrast, generating entire 3D scenes along a single, large camera path remains challenging due to several reasons. First, defining a continuous path a-priori, that covers an entire 3D scene is ambiguous. In other words, we have to adapt the trajectory on-the-fly based on the generated content (e.g., prevent running into walls, resolve occlusions, discover novel generated areas). Second, to achieve 3D-consistency in large 3D worlds, the model must memorize previously generated objects/structures across a large history of previous frames. While approaches like  (Zhou et al., 2025; Song et al., 2025) show promising results, their models still suffer from inconsistencies, collisions, and catastrophic forgetting, when re-visiting previously observed areas (also see Figure 4 and the supplementary material).
@@
Crucially, resolving collisions becomes a much easier task this way: since we already split scene generation into discontinuous and small camera paths, we only need to prevent such degenerate cases for each trajectory separately. In other words, the next trajectory is not impacted by a collision in previous trajectories: it can still be generated starting from the scene scaffold. Concretely, we first generate the entire video and then utilize a video depth estimator (Chen et al., 2025a) to predict normalized depth maps $\mathbf{\hat{D}}\in[0,1]^{N\times H\times W}$ for all generated frames. We define a fixed area of interest as a binary mask $\mathbf{M}\in\{0,1\}^{H\times W}$, where pixels $(u,v)$ within a center crop of extent $(w,h)$ are set to 1:
@@
Our method leverages the two-stage video generation approach of (Zhou et al., 2025), where anchor frames are first generated uniformly along the trajectory, followed by interpolation between these anchors. Our scene memory mechanism ensures consistency of the anchor frames across trajectories. The subsequent interpolation then naturally produces scene-consistent intermediate frames.